Fermi National Laboratory
 

Accelerator Update for July 21, 2006 - July 24, 2006

 

Friday July 21

The day shift began with Operations monitoring store 4845 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with the NuMI horn cooling work continuing, and with MiniBooNE taking beam.

Operations prepared the Recycler at 9:30 AM to stash antiprotons. The stash ended and Pbar resumed stacking at 10:11 AM.

Operations established beam to the Meson MTest experiment at 12:15 PM.

At 12:57 PM, Operations reported that a Linac RF station (LRF3) had tripped off for the first of four times in about 30 minutes. It reset ever time except for the last. Linac experts reported at 1:26 PM that the station was sparking and that they were taking it down to repair.

At 1:48 PM, the H- Source alarmed; two vacuum pumps had tripped off. An expert reset the pump.

The Recycler began preparing to stash at 2:02 PM.

Operations terminated store 4845 at 2:10 PM.

Linac experts reported at 2:26 PM that they had completed their repairs on LRF3. Beam resumed.

Operations prepared CDF and D0 for an access at 2:27 PM.

At 2:35 PM, the H- Source vacuum pumps tripped off again. Beam resumed ten minutes later.

The stash ended and Pbar resumed stacking at 3:14 PM. MTest resumed taking beam at 3:37 PM.

CDF came out of access at 3:48 PM.

D0 came out of access at 4:15 PM.

Operations put the TeV into a dry squeeze at 4:42 PM.

Operations put the TeV into shot setup at 5:05 PM.

At 6:02 PM, Operations disabled the MTest beamline and dropped their interlocks after discovering there had been a Controlled Access violation. MTest will get no beam until an investigation has been completed and corrective measures have been implemented.

Operations established store 4847 at 6:48 PM with an initial luminosity of 149.34E30.

At 7:20 PM, LRF5 tripped off. Operators and an expert responded. Pbar and MiniBooNE resumed taking beam by 8 PM.

Saturday July 22

The midnight shift began with Operations monitoring store 4847, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler began preparing to stash at 1:29 AM.

At 1:58 AM, the Booster West Anode Power Supply (WAPS) tripped off for the first of two times in about forty-five minutes. It reset both times.

The stash ended at 2:20 AM.

At 15:56 AM, WAPS tripped off for the first of two times in fifteen minutes. It reset both times.

The Recycler completed a stash and Pbar resumed stacking at 6:26 AM.

At 8:43 AM, MI aborted. Operators had trouble resetting due to a ground fault indication. After talking with experts, operators determined that there was a MI-40 quadrupole power supply problem. A power supply expert came in and replaced a bad SCR.

Pbar resumed stacking at 11:06 AM.

At 11:23 AM, Operations reported that the Booster was having some beam position problems. A Booster expert discovered problems with BRF17 & 18. RF experts came in to investigate this and a MIRF problem.

At 12:11 PM, an RF expert reported that MIRF4 had a bias grid short and would require a tunnel access to repair.

RF experts brought up BRF17 & 18 at 12:40 PM. They reported that station BRF18 had a bad power supply.

The Recycler began preparing to stash at 1:11 PM. Pbar resumed stacking at 1:46 PM.

The Recycler began preparing to stash at 5:07 PM. Pbar resumed stacking at 5:34 PM.

Operations reported at 8:52 PM that they were in the middle of a stash. The stash ended and Pbar resumed stacking at 9:16 PM.

Sunday July 23

The midnight shift began with Operations monitoring store 4847 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler began preparing to stash at 12:56 AM. The stash ended at 1:40 AM. Pbar resumed stacking.

Operations terminated store 4847 at 4:01 AM.

Operations began preparing D0 for an access at 4:09 AM. The access ended at 5:27 AM.

Operations put the TeV through a dry squeeze at 5:48 AM and then put the TeV into shot setup.

The Recycler began preparing to stash at 8:27 AM.

Operations established store 4851 at 8:31 AM with an initial luminosity of 168.32E30.

The stash ended and Pbar resumed stacking at 9:06 AM.

The Recycler began preparing to stash at 12:31 PM. The stash ended and Pbar resumed stacking at 1:19 PM.

The Recycler began preparing to stash at 5:06 PM. The stash ended and Pbar resumed stacking at 5:33 PM.

At 6:23 PM, an Accumulator RF station (ARF1) tripped off due to a CAMAC crate power supply breaker trip. Operators reported at 7:24 PM that they had reset the breaker, but that little beam was making it into the Accumulator. Operators restored a crate file and this improved the stacking rate a little. Operations called in Pbar experts.

Pbar experts checked out ARF1 at 9:48 PM. Later they reported that they had cleaned the connections for the ARF1 feedback system. They also reported a frequency offset that will be investigated further on Monday.

The Recycler began preparing to stash at 10:31 PM. The stash ended and Pbar resumed stacking at 11:03 PM.

Monday July 24

The midnight shift began with Operations monitoring store 4851 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 3:07 AM, Operations reported problems with the Recycler’s electron cooling.

The Recycler began preparing to stash at 3:31 AM. The stash ended and Pbar resumed stacking at 3:51 AM.

The Recycler began preparing to stash at 7:20 AM.

The Plans for Monday

The plans for today are to stash, stack, and store until this afternoon. Operations will establish another store and then allow Pbar to make a 16-hour access for repairs.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/24/2006    email Fermilab